# Category-free Out-of-Distribution Node Detection with Feature Resonance

### v1

@@ Abstract
Detecting out-of-distribution (OOD) nodes in the graph-based machine-learning field is challenging, particularly when in-distribution (ID) node multi-category labels are unavailable. Thus, we focus on feature space rather than label space and find that, ideally, during the optimization of known ID samples, unknown ID samples undergo more significant representation changes than OOD samples, even if the model is trained to fit random targets, which we called the Feature Resonance phenomenon. The rationale behind it is that even without gold labels, the local manifold may still exhibit smooth resonance. Based on this, we further develop a novel graph OOD framework, dubbed Resonance-based Separation and Learning (RSL), which comprises two core modules: (i)-a more practical micro-level proxy of feature resonance that measures the movement of feature vectors in one training step. (ii)-integrate with synthetic OOD nodes strategy to train an effective OOD classifier. Theoretically, we derive an error bound showing the superior separability of OOD nodes during the resonance period. Empirically, RSL achieves state-of-the-art performance, reducing the FPR95 metric by an average of 18.51% across five real-world datasets.
Machine Learning, ICML
@@ 1 Introduction
(a) Toy Dataset
(b) Gradient Descent Trajectory
Figure 1: (a) We conduct a preliminary study on the changes in ID and OOD node representations during training using a toy dataset. (b) Projections of the representations of ID and OOD nodes onto gradients: $\text{Proj}_{\nabla\ell(\theta_{t};\cdot)}\mathbf{x}_{i}=\frac{\mathbf{x}_{i}\cdot\nabla\ell(\theta_{t};\cdot)}{\parallel\nabla\ell(\theta_{t};\cdot)\parallel_{2}^{2}}\cdot\nabla\ell(\theta_{t};\cdot)$.
@@
Graph-based machine learning models like Graph Neural Networks (GNNs) (Kipf & Welling, 2016a; Xu et al., 2018; Abu-El-Haija et al., 2019) have become increasingly prevalent in applications such as social network analysis (Fan et al., 2019), knowledge graphs (Baek et al., 2020), and biological networks (De Cao & Kipf, 2018). Despite the success of GNNs, detecting out-of-distribution (OOD) nodes remains an under-explored challenge. These OOD nodes differ significantly from the in-distribution (ID) nodes used during training, and their presence can severely undermine the performance and robustness of graph models. As deploying GNNs in real-world environments becomes more common, the ability to identify and handle OOD nodes is crucial for ensuring the reliability of using these models.
To address this, most existing methods (Hendrycks & Gimpel, 2016; Liang et al., 2017; Hendrycks et al., 2018; Liu et al., 2020; Wu et al., 2023) assume that ID nodes are equipped with multi-category labels. Then, they train an in-distribution classifier and develop OOD metrics based on (i)-classifier outputs, such as Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016) and Energy (Liu et al., 2020; Wu et al., 2023); (ii)-supervised representations, such as KNN (Sun et al., 2022) and NNGuide (Park et al., 2023). However, in many real-world scenarios, accessing multi-category labels for ID nodes is often highly challenging due to practical limitations such as the high cost of annotation, lack of domain expertise, or data privacy concerns, which essentially hinder the effectiveness of current OOD methods. To date, only a few papers (Gong & Sun, 2024; Sun et al., 2022; Sehwag et al., 2021; Liu et al., 2023) study this practical setup, and there is still a large room for improvement, especially in the graph field at the node level.
@@
In this paper, we revisit the graph OOD task at the node level from a new perspective and turn our attention to the intrinsic similarities within the data. An intuitive idea is that the ID samples may still share some commonalities in the representation space. We hypothesize that when optimizing the representation of known ID nodes, the representation of unknown ID nodes and unknown OOD nodes will change with different trajectories. Based on the hypothesis and using a toy dataset (Figure 1(a)), we design an experiment where the features of labeled ID samples are aligned to an arbitrarily fixed representation vector. Interestingly, we observe a distinct behavior during this optimization process: the representations of unlabeled wild ID samples experienced more pronounced changes than wild OOD samples, as shown in Figure 1(b). This phenomenon closely resembles the concept of forced vibration, where resonance occurs when an external force aligns with the natural frequency of an oscillator, amplifying its oscillation to a maximum. Analogously, we refer to this phenomenon as Feature Resonance: during the optimization of known ID samples, the representation of unknown ID samples undergoes more significant changes compared to OOD samples. This phenomenon reveals the intrinsic relationship between ID samples, highlighting their shared underlying distribution. Evidently, this feature resonance phenomenon can be leveraged for OOD detection: weaker representation changes during known ID optimization indicate a higher likelihood of being OOD.
@@
In real-world scenarios, due to the intrinsic complex pattern in data, we find that the feature resonance phenomenon still occurs but slightly differs from the ideal conditions. To illustrate this, we further propose a micro-level proxy for measuring feature resonance—by computing the movement of the representation vector in one training step. Our findings reveal that in more complex scenarios, the feature resonance phenomenon typically arises during the middle stages of the training process, whereas during other phases, it may be overwhelmed by noise or obscured by overfitting. In such cases, evaluating the entire trajectory often fails to yield satisfactory results. Fortunately, efficient OOD detection can still be achieved by calculating the micro-level feature resonance measure. By utilizing a simple binary ID/OOD validation set [^1] The use of the validation set is consistent with previous works (Katz-Samuels et al., 2022; Gong & Sun, 2024; Du et al., 2024a, b) and does not contain multi-category labels., we empirically show the feature resonance period can be precisely identified, and we identify more minor representation differences as OOD samples. Notably, our new micro-level feature resonance measure is still label-independent by fitting a randomly fixed target, making it highly compelling in category-free scenarios. Theoretical and experimental proof that micro-level feature resonance can filter a set of reliable OOD nodes with low error, e.g., on the FPR95 metric, the micro-level feature resonance achieves an average reduction of 10.93% compared to current state-of-the-art methods.
Furthermore, we combine the micro-level feature resonance with the current Langevin-based synthetic OOD nodes generating strategy to train an OOD classifier for more effective OOD node detection performance, which we call the whole framework as RSL; for example, the FPR95 metric is reduced by an average of 18.51% compared to the current state-of-the-art methods.
2 Preliminaries
Graph Neural Network. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ denote an graph, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of nodes and edges, respectively. We represent the node space by $\mathbb{V}$, and $\mathcal{V}\in\mathbb{V}$. ${\boldsymbol{X}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ denote the matrix of features of the nodes. Here, the representation of a node $v$ can be defined as $\mathbf{h}_{v}$. Graph neural networks (GNNs) aim to update the representation of the given graph $\mathcal{G}$ by leveraging its topological structure. For the representation $\mathbf{h}_{v}$ of node $v$, its propagation of the $k$-th layer GNN is represented as:
$\begin{array}[]{cl}\mathbf{h}_{v}^{(k)}&=g(\mathbf{h}_{v}^{(k-1)};\theta)\\ &=\mathrm{UP}^{(k)}\big{(}\mathrm{AGG}^{(k)}\big{(}\mathbf{h}_{u}^{(k-1)}:\forall u\in\mathcal{N}(v)\cup v\big{)}\big{)}\end{array}$ (1)
where $g(\cdot;\theta)$ denotes the GNN encoder, $\theta$ represents all trainable parameters of GNN encoder. $\mathrm{AGG}(\cdot)$ denotes a function that aggregates messages from the neighbors of node $v$, $\mathcal{N}(v)$ represents the set of neighbors. $\mathrm{UP}(\cdot)$ denotes a function that updates the representation of node $v$ with the current representation of $v$ and the aggregated vector.
Problem Statement. The node set can be divided into $\mathcal{V}=\mathcal{V}_{\mathrm{in}}\cup\mathcal{V}_{\mathrm{out}}$, where $\mathcal{V}_{\mathrm{in}}$ and $\mathcal{V}_{\mathrm{out}}$ represents the ID node set and OOD node set, respectively. We assume ID nodes are sampled from the distribution $\mathbb{P}_{\mathrm{in}}$, and OOD nodes are sampled from distribution $\mathbb{P}_{\mathrm{out}}$. We formally define the Category-free OOD node detection task:
Definition 2.1.
Category-free OOD node detection. Given a collection of nodes sampled from $\mathbb{P}_{\mathrm{in}}$ and $\mathbb{P}_{\mathrm{out}}$, the objective is to correctly identify the source of each node, whether it is from the $\mathbb{P}_{\mathrm{in}}$ or $\mathbb{P}_{\mathrm{out}}$.
Unlabeled Wild Node. In this work, we incorporate unlabeled wild node $\mathcal{V}_{\mathrm{wild}}=\{\tilde{v}_{1},\cdots,\tilde{v}_{m}\}$ with feature $\boldsymbol{X}_{\mathrm{wild}}=\{\tilde{\mathbf{x}}_{1},\cdots,\tilde{\mathbf{x}}_{m}\}$ into our learning framework, leveraging the fact that such features are often accessible. We define the unlabeled wild nodes distribution as follows:
Definition 2.2.
Unlabeled wild nodes. Unlabeled wild nodes typically consist of a mixture of ID nodes and OOD nodes. We use the Huber contamination model (Huber, 1992) to characterize the marginal distribution of the wild data:
$\mathbb{P}_{\mathrm{wild}}=(1-\pi)\mathbb{P}_{\mathrm{in}}+\pi\mathbb{P}_{\mathrm{out}}$ (2)
where $\pi\in(0,1]$.
Figure 2: Schematic of Feature Resonance.
@@ 3.1 Revealing the Feature Resonance Phenomenon
Previous studies (Hendrycks & Gimpel, 2016; Liu et al., 2020; Wu et al., 2023) mostly train a classifier on ID nodes with multi-category labels and develop selection criteria based on output probabilities, e.g. entropy. However, these methods become inapplicable in category-free scenarios.
@@
[…]
Despite the promise, the abovementioned feature resonance phenomenon occurs under multi-category training. But how can we induce this phenomenon in a category-free scenario without multi-category labels? Interestingly, we find that even when random labels are assigned to known ID nodes for multi-category training, the trajectories of unknown ID nodes are still more significant than those of unknown OOD nodes. More surprisingly, on a ideal toy dataset, even when all known ID node representations are aligned toward one single random fixed target vector, the trajectories of unknown ID nodes are still longer than those of unknown OOD nodes, as shown in Figure 1. Green points represent unknown ID samples, blue points represent unknown OOD samples, and red points represent known ID samples aligned to a target vector. As shown in Figure 1(b), modifying the representation of known ID samples results in longer representation change trajectories for unknown ID samples compared to unknown OOD samples. The experiments above indicate that the feature resonance phenomenon is label-independent and results from the intrinsic relationships between ID node representations. Therefore, this is highly suitable for category-free OOD detection scenarios without multi-category labels.
@@
Definition 3.1.
Feature Resonance (macroscopic): For any optimization objective $\ell(\boldsymbol{X}_{\text{known}},\cdot)$ applied to the representations $\boldsymbol{X}_{\text{known}}$ of known ID samples derived from any model $h_{\theta}(\cdot)$, we have $\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$.
@@ 3.2 Utilizing the Micro-level Feature Resonance Phenomenon with An Arbitrary Target
$\ell(h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{1}^{\top}e-({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (4)
where $h_{\theta_{t}}({\boldsymbol{X}}_{\text{known}})={\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top}$ represent the last linear layer of the model for representation transformation and $e$ denotes an arbitrary randomly generated target vector.
@@
Definition 3.2.
Feature Resonance (microscopic): For any optimization objective $\ell(\boldsymbol{X}_{\text{known}},\cdot)$ applied to the known ID nodes’ representations $\boldsymbol{X}_{\text{known}}$ from any model $h_{\theta_{t}}(\cdot)$, during the optimization process, there exists $t$ such that $\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$. We define the resonance-based filtering score as $\tau_{i}=\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{2}$. The resonance-based scores $\tau$ of OOD nodes should be smaller than those of ID nodes at $t$.
By observing $\tau$ for ID samples and OOD samples, we find that feature resonance does not persist throughout the entire training process but rather occurs at specific stages of training. In our experiments on the common benchmarks, we find that during the early stages of training, the model is searching for the optimal optimization path, leading to chaotic representation changes and thus making feature resonance insignificant. However, in the middle stages of training, once the model identifies an optimization path that aligns with the patterns of the ID samples, it optimizes along the path most relevant to the features of the ID samples, and feature resonance becomes most prominent. As the model continues to optimize and enters the overfitting stage, the feature resonance phenomenon begins to dissipate. Figure 3 shows the experimental results on the Amazon dataset, and others are provided in Figure 6 of the Appendix. Through the above experiments and analyses, we find that using $\hat{F}(\tilde{\mathbf{x}}_{i})$ to identify OOD nodes is affected by error accumulation and is, therefore, not a reliable approach. However, there exists a specific period during training when micro-level feature resonance occurs. By utilizing a validation set (Katz-Samuels et al., 2022; Gong & Sun, 2024; Du et al., 2024a, b), we can easily identify the period during which feature resonance occurs.
@@
$\begin{split}&g_{\gamma}(\tilde{\mathbf{x}}_{i})=\mathds{1}\{\tau_{i}^{*}\leq\gamma\},\\ &\text{s.t.},\tau^{*}=\max_{t}\text{AUROC}(\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{in}}},\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{out}}})\end{split}$ (5)

### v2

# Harnessing Feature Resonance under Arbitrary Target Alignment for Out-of-Distribution Node Detection

@@ Abstract
Out-of-distribution (OOD) node detection in graphs is a critical yet challenging task. Most existing approaches rely heavily on fine-grained labeled data to obtain a pre-trained supervised classifier, inherently assuming the existence of a well-defined pretext classification task. However, when such a task is ill-defined or absent, their applicability becomes severely limited. To overcome this limitation, there is an urgent need to propose a more scalable OOD detection method that is independent of both pretext tasks and label supervision. We harness a new phenomenon called Feature Resonance, focusing on the feature space rather than the label space. We observe that, ideally, during the optimization of known ID samples, unknown ID samples undergo more significant representation changes than OOD samples, even when the model is trained to align arbitrary targets. The rationale behind it is that even without gold labels, the local manifold may still exhibit smooth resonance. Based on this, we further develop a novel graph OOD framework, dubbed Resonance-based Separation and Learning (RSL), which comprises two core modules: (i)-a more practical micro-level proxy of feature resonance that measures the movement of feature vectors in one training step. (ii)-integrate with a synthetic OOD node strategy to train an effective OOD classifier. Theoretically, we derive an error bound showing the superior separability of OOD nodes during the resonance period. Extensive experiments on a total of thirteen real-world graph datasets empirically demonstrate that RSL achieves state-of-the-art performance. The code is available via https://github.com/ShenzhiYang2000/RSL.
@@ 1 Introduction
Graph-based machine learning models like Graph Neural Networks (GNNs) (Kipf and Welling, 2016a; Xu et al., 2018; Abu-El-Haija et al., 2019; Zhou et al., 2024) have become increasingly prevalent in applications such as social network analysis (Fan et al., 2019), knowledge graphs (Baek et al., 2020), and biological networks (De Cao and Kipf, 2018). Despite the success of GNNs, detecting out-of-distribution (OOD) nodes remains an under-explored challenge. These OOD nodes differ significantly from the in-distribution (ID) nodes used during training, and their presence can severely undermine the performance and robustness of graph models. As deploying GNNs in real-world environments becomes more common, the ability to identify and handle OOD nodes is crucial for ensuring the reliability of using these models.
To address this, most existing methods (Hendrycks and Gimpel, 2016; Liang et al., 2017; Hendrycks et al., 2018; Liu et al., 2020; Wu et al., 2023) employ classifiers pretrained on a preceding classification task to develop OOD metrics based on (i)-classifier outputs, such as Maximum Softmax Probability (MSP) (Hendrycks and Gimpel, 2016) and Energy (Liu et al., 2020; Wu et al., 2023); (ii)-supervised representations, such as KNN (Sun et al., 2022) and NNGuide (Park et al., 2023). These methods heavily rely on two key assumptions: (i) the availability of multi-class labels, and (ii) a well-defined pretext multi-class classification task. However, in practice, there exists a wide range of OOD detection scenarios that fall outside these constraints. In many cases, the pretext task is not classification—for example, OOD detection in generative modeling (Nalisnick et al., 2018; Ren et al., 2019), regression (Lakshminarayanan et al., 2017), or reinforcement learning (Nasvytis et al., 2024). In some scenarios, there may not even be a defined pretext task at all, such as in one-class OOD detection (Ruff et al., 2018). These non-classification settings lack accessible multi-class labels, making it difficult to directly apply existing methods. Therefore, there is an urgent need for label-agnostic and unsupervised approaches that can operate effectively in such contexts. To date, only a few papers (Gong and Sun, 2024; Sehwag et al., 2021; Liu et al., 2023) study this practical setup, and there is still a large room for improvement, especially in the graph field at the node level.
(a) (a) We conduct a preliminary study on the changes in ID and OOD node representations during training using a toy dataset. (b) Projections of the representations of ID and OOD nodes onto gradients: $\text{Proj}_{\nabla\ell(\theta_{t};\cdot)}\mathbf{x}_{i}=\frac{\mathbf{x}_{i}\cdot\nabla\ell(\theta_{t};\cdot)}{\parallel\nabla\ell(\theta_{t};\cdot)\parallel_{2}^{2}}\cdot\nabla\ell(\theta_{t};\cdot)$. (c) Schematic of Feature Resonance.
@@
In this paper, we revisit the graph OOD detection task at the node level from a new perspective and turn our attention to the intrinsic similarities within the data. An intuitive idea is that the ID samples may still share some commonalities in the representation space. We hypothesize that when optimizing the representation of known ID nodes, the representation of unknown ID nodes and unknown OOD nodes will change with different trajectories. Based on the hypothesis and using a toy dataset (Figure 3(a)(a)), we design an experiment where the features of labeled ID samples are aligned to an arbitrarily fixed representation vector. Interestingly, we observe a distinct behavior during this optimization process: the representations of unlabeled wild ID samples experienced more pronounced changes than wild OOD samples, as shown in Figure 3(a)(b). This phenomenon closely resembles the concept of forced vibration, where resonance occurs when an external force aligns with the natural frequency of an oscillator, amplifying its oscillation to a maximum. Analogously, we refer to this phenomenon as Feature Resonance: during the optimization of known ID samples, the representation of unknown ID samples undergoes more significant changes compared to OOD samples. This phenomenon reveals the intrinsic relationship between ID samples, highlighting their shared underlying distribution. Evidently, this feature resonance phenomenon can be leveraged for OOD detection: weaker representation changes during known ID optimization indicate a higher likelihood of being OOD.
@@
In real-world scenarios, due to the intrinsic complex pattern in data, we find that the feature resonance phenomenon still occurs but slightly differs from the ideal conditions. To illustrate this, we further propose a micro-level proxy for measuring feature resonance—by computing the movement of the representation vector in one training step. Our findings reveal that in more complex scenarios, the feature resonance phenomenon typically arises during the middle stages of the training process, whereas during other phases, it may be overwhelmed by noise or obscured by overfitting. In such cases, evaluating the entire trajectory often fails to yield satisfactory results. Fortunately, efficient OOD detection can still be achieved by calculating the micro-level feature resonance measure. By utilizing a simple binary ID/OOD validation set [^1] The use of the validation set is consistent with previous works (Katz-Samuels et al., 2022; Gong and Sun, 2024; Du et al., 2024a, b) and does not contain multi-category labels., we empirically show the feature resonance period can be precisely identified, and we identify more minor representation differences as OOD samples. Notably, our new micro-level feature resonance measure is still label-independent by fitting a randomly fixed target, making it highly compelling in category-free and task-agnostic scenarios. Theoretical and experimental proof that micro-level feature resonance can filter a set of reliable OOD nodes with low error. Furthermore, we combine the micro-level feature resonance with the current Langevin-based synthetic OOD nodes generating strategy to train an OOD classifier for more effective OOD node detection performance, which we call the whole framework as RSL; for example, the FPR95 metric is reduced by an average of 15.20% compared to the current state-of-the-art methods.
@@ 2.1 Revealing the Feature Resonance Phenomenon
Previous studies (Hendrycks and Gimpel, 2016; Liu et al., 2020; Wu et al., 2023) mostly train a classifier on ID nodes with multi-category labels and develop selection criteria based on output probabilities, e.g., entropy. However, these methods become inapplicable in category-free and task-agnostic scenarios.
@@
In our preliminary experiments, we first calculate the metric under supervised conditions and observe a significant difference between the feature trajectories of ID samples and those of OOD samples. Specifically, we perform multi-category training on known ID nodes on two datasets with true $N$-category labels, Squirrel and WikiCS  [^2] NN is the number of categories, and the results above with different target vectors are shown in Table 6.. Imagine that during multi-category training, representations of known ID nodes within the same category align while unknown ID nodes drift toward the corresponding category centers. However, the trajectory trends and lengths of unknown ID nodes differ significantly from those of OOD nodes, with the former showing more distinct trends and longer trajectories; see Figure 3(a) (c) for visual illustration. In other words, the well-defined in-distribution manifold is always shaped by ID samples, whose representation trajectories tend to exhibit similar behavior, which we refer to as feature resonance. Conversely, OOD samples belong to distinct manifold structures, making their representations less likely to converge coherently. Evidently, this feature resonance phenomenon can be leveraged for OOD detection.
Despite the promise, the abovementioned feature resonance phenomenon occurs under multi-category training. But how can we induce this phenomenon in a label-agnostic scenario without multi-category labels? Interestingly, we find that even when random labels are assigned to known ID nodes for multi-category training, the trajectories of unknown ID nodes are still more significant than those of unknown OOD nodes. More surprisingly, on a ideal toy dataset, even when all known ID node representations are aligned toward one single random fixed target vector, the trajectories of unknown ID nodes are still longer than those of unknown OOD nodes, as shown in Figure 3(a)(a). Green points represent unknown ID samples, blue points represent unknown OOD samples, and red points represent known ID samples aligned to a target vector. As shown in Figure 3(a)(b), modifying the representation of known ID samples results in longer representation change trajectories for unknown ID samples compared to unknown OOD samples. The experiments above indicate that the feature resonance phenomenon is label-independent and results from the intrinsic relationships between ID node representations. Therefore, this is highly suitable for category-free and task-agnostic OOD detection scenarios without multi-category labels.
@@
Definition 1.
Feature Resonance (macroscopic): For any optimization objective $\ell(\bm{X}_{\text{known}},\cdot)$ applied to the representations $\bm{X}_{\text{known}}$ of known ID samples derived from any model $h_{\theta}(\cdot)$, we have $\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\hat{F}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$.
@@ 2.2 Utilizing the Micro-level Feature Resonance Phenomenon with An Arbitrary Target
$\ell(h_{\theta_{t}}({\bm{X}}_{\text{known}}),e)=\mathbb{E}(\parallel\mathbf{1}^{\top}e-({\bm{X}}_{\text{known}}\mathbf{W}^{\top})\parallel^{2}_{2})$ (2)
where $h_{\theta_{t}}({\bm{X}}_{\text{known}})={\bm{X}}_{\text{known}}\mathbf{W}^{\top}$ represent the last linear layer of the model for representation transformation and $e$ denotes an arbitrary randomly generated target vector.
@@
Definition 2.
Feature Resonance (microscopic): For any optimization objective $\ell(\bm{X}_{\text{known}},\cdot)$ applied to the known ID nodes’ representations $\bm{X}_{\text{known}}$ from any model $h_{\theta_{t}}(\cdot)$, during the optimization process, there exists $t$ such that $\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{in}}}>\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{\mathbb{P}^{\mathrm{wild}}_{\mathrm{out}}}$. We define the resonance-based filtering score as $\tau_{i}=\parallel\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\parallel_{2}$. The resonance-based scores $\tau$ of OOD nodes should be smaller than those of ID nodes at $t$.
By observing $\tau$ for ID samples and OOD samples, we find that feature resonance does not persist throughout the entire training process but rather occurs at specific stages of training. In our experiments on the common benchmarks, we find that during the early stages of training, the model is searching for the optimal optimization path, leading to chaotic representation changes and thus making feature resonance insignificant. However, in the middle stages of training, once the model identifies an optimization path that aligns with the patterns of the ID samples, it optimizes along the path most relevant to the features of the ID samples, and feature resonance becomes most prominent. As the model continues to optimize and enters the overfitting stage, the feature resonance phenomenon begins to dissipate. Figure 3 shows the experimental results on the Amazon dataset, and others are provided in Figure 7(a) of the Appendix. Through the above experiments and analyses, we find that using $\hat{F}(\tilde{\mathbf{x}}_{i})$ to identify OOD nodes is affected by error accumulation and is, therefore, not a reliable approach. However, there exists a specific period during training when micro-level feature resonance occurs. By utilizing a validation set (Katz-Samuels et al., 2022; Gong and Sun, 2024; Du et al., 2024a, b), we can easily identify the period during which feature resonance occurs.
@@
$\begin{split}&g_{\gamma}(\tilde{\mathbf{x}}_{i})=\mathds{1}\{\tau_{i}^{*}\leq\gamma\},\ \ \ \text{s.t.},\tau^{*}=\max_{t}\text{AUROC}(\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{in}}},\tau^{t}_{\mathcal{V}_{\mathrm{val}}^{\mathrm{out}}})\end{split}$ (3)